# thanks: Equal contributionsthanks: Equal contributions

### v2

@@ Abstract
Edge states in topological systems have attracted great interest due to their robustness and linear dispersions. Here a superconducting-proximitized edge interferometer is engineered on a topological insulator Ta2Pd3Te5 with asymmetric edges to realize the interfering Josephson diode effect (JDE), which hosts many advantages, such as the high efficiency as much as 73% at tiny applied magnetic fields with an ultra-low switching power around picowatt, and a giant interfering magnetochiral anisotropy with a maximal coefficient $\gamma$ = $1.2\times 10^{9}$ T-1A-1. As an important element to induce such JDE, the second-order harmonic in the current-phase relation is also experimentally confirmed by half-integer Shapiro steps. This edge interferometer offers a novel and effective method to enhance the overall performance of JDE and magnetochiral anisotropy, and boosts great potential applications for future superconducting quantum devices.
@@
Figure 1: Mechanisms for SDE/JDE. $\mathbf{a}$, JDE in the asymmetric SQUID. $\mathbf{b}$, An asymmetric SQUID formed by asymmetric edge states in Ta2Pd3Te5 JJ and its device configuration. The upper and lower edges are the palladium and tantalum-tellurium atomic chain, respectively. The inset shows the photomicrograph of device S2, and the white scale bar corresponds to 1 $\mu$m. $\mathbf{c-f}$, Simulated CPR based on the minimal model. JDE only exists in $\mathbf{d}$ with different supercurrents, a non-zero magnetic field ($\phi/\phi_{0}\neq 0$) and higher harmonics in the CPR ($\alpha_{n}\neq 0$). $\mathbf{g}$, Band structure of monolayer Ta2Pd3Te5. Edge states (blue and red lines) exist near the Fermi level and are marked in the corresponding edges of $\mathbf{b}$. $\mathbf{h}$, $\left|I_{b}\right|-\left|V\right|$ curves for positive and negative current sweep at $B_{z}$ = 8.4 mT. $\mathbf{i,j}$, Alternating switching between the superconducting and normal states at $B_{z}$ = 8.4 mT and 10 mK in device S2.
@@
Essentially, the appearance of JDE is related to the difference between maximum($I_{c+}$) and minimum($I_{c-}$) current in the CPR. The conventional CPR is an odd function of superconducting phase difference $\varphi$ between superconducting leads, entailing a zero supercurrent at $\varphi$ = 0, which shifts [meaning $I(\varphi=0)$neq 0] when both time-reversal and chiral symmetries are broken 30; 21; 31. To achieve JDE, other mechanisms to cause the CPR deviating from standard sinusoidal $I(\varphi)$ = $I_{c}$ $sin(\varphi)$ will also be introduced. The characteristics of CPR, which are usually inferred by the interference pattern of SQUIDs and the Shapiro steps due to ac Josephson effect, as well as the newly-discovered JDE with non-reciprocal critical current, have been widely harnessed to detect the anomalous superconductivity 32; 33, such as proximitized helical/chiral topological edge states. In turn, in a SQUID constructed by two asymmetric edge states of the high-order topological insulator (HOTI), JDE with an unconventional CPR could be easily realized with highly-tunable efficiency 5 and relatively easy fabrication (Fig. 1b). The relatively small number of edge supercurrent channels, in comparison to the case of bulk states transport, may contribute to the small critical current and power consumption 33; 34.
@@
More precisely, the interferometer formed by two asymmetric edge supercurrents (Fig. 1b) can be viewed as a SQUID with two different JJs, and induces JDE with requirements similar to the asymmetric SQUID 9 in Fig. 1a. According to the Souto $et$ $al$. 9, a minimal model with $n$ JJs concatenated in an interferometer array is applied to account for the origin of JDE, and the CPR only takes the first and second harmonic contribution into consideration. The current is written as $I(\varphi,\phi/\phi_{0})=\sum_{1}^{n=2}I_{n}sin(\varphi+2\pi(n-1)\phi/\phi_{0})+\alpha_{n}I_{n}sin(2\varphi+4\pi(n-1)\phi/\phi_{0})$ for two JJs in parallel, where $I_{n}$ ($\alpha_{n}I_{n}$) is the amplitude of the first (second) harmonic content for the $n$th JJ, $\alpha_{n}$ is current coefficient of the higher harmonic, $\phi$ is the magnetic flux and $\phi_{0}=h/2e$ is the magnetic flux quantum. In Fig. 1c-f, the simulations of the CPR using the aforementioned formula serves the purpose of illustrating three key requirements that contribute to the difference between $I_{c+}$ (red arrow) and $I_{c-}$ (pink arrow). First, an external magentic field ($\phi/\phi_{0}\neq 0$) is necessary to break time-reversal symmetry and induce nonzero current at $\varphi$ = 0 in the CPR (Fig. 1d-f). Second, the supercurrent of each edge, including all the harmonic contributions, should be different (see Fig. 1d,e and detailed simulations in supplementary Fig. S2). This condition can be realized by two interfering edge states with different dispersion 5. It can also be met, for example, by two superconducting quantum point contacts with different transmission coefficients 9, which may be linked to various factors such as different edge states, disorder situations, among others. All of these can cause the different supercurrent of edges in our devices, which is actually quite unique in topological materials. Finally, at least one edge channel should be transmissive so that the CPR acquires a higher harmonic ($\alpha_{n}\neq 0$) 9; 35 (see Fig. 1d, f).
@@
In this work, highly efficient JDE, residing in the tilting supercurrent pattern, is reported in a Ta2Pd3Te5 edge interferometer at very small $B_{z}$ and is driven by low power consumption to achieve a highly stable rectification effect. It is also accompanied with fractional Shapiro steps, revealing the higher harmonic contributions of the non-negligible transmission channels and predominantly enhanced interfering magnetochiral anisotropy (MCA). The choice of Ta2Pd3Te5 is based on the following reasons: (1) Ta2Pd3Te5 is a van der Waals material with quasi-one-dimensional (1D) chains 36 and can be easily mechanically exfoliated. (2) As a 2D topological insulator, where multiple layers could be viewed as simple stacking of monolayers due to very weak inter-layer coupling 37; 38, it hosts excitonic insulator states 39; 40; 41 and edge states observed by scanning tunnelling microscopy 42; 40 and transport measurements 38. Notably, owing to the anisotropic bonding energy of 1D chains, the edges possess different atomic chains (Pd and Ta-Te atomic chains), resulting in asymmetric dispersion, as shown by the red and blue lines/arrows in Fig. 1b, g. (3) Superconductivity can be easily achieved in this system through doping or high-pressure techniques 43; 44; 45. Therefore, by proximity effect, the induced supercurrent (supplementary Fig. S1) carried by edge states of Ta2Pd3Te5 can be used to realize an asymmetric SQUID with the interfering JDE.
Figure 2: Asymmetric Josephson effect and JDE in Ta2Pd3Te5 JJs. $\mathbf{a}$, SQUID pattern of device S1 at 15 mK. $\mathbf{b}$, Position-dependent supercurrent density distribution. Inset: A sketch of edge-supercurrents-formed JJ. $\mathbf{c}$, $B_{z}$-dependent $I_{c+}$ and $\left|I_{c-}\right|$. $\mathbf{d}$, Non-reciprocal critical current $\Delta I_{c}=I_{c+}-\left|I_{c-}\right|$, Josephson diode efficiency $\eta$ and fitted $\Delta I_{c}$ (red line) as a function of $B_{z}$. $\mathbf{e}$, Temperature-dependent $I_{c+}$, $I_{c-}$, $\Delta I_{c}$ and $\eta$ for device S1 at $B_{z}$ = 6.4 mT. $\mathbf{f}$, SQUID pattern of device S2 at 10 mK. $\mathbf{g}$, Supercurrent density distribution. The inset is a draft of the JJ with three edge supercurrents. $\mathbf{h}$, $B_{z}$ dependence of $I_{c+}$ and $\left|I_{c-}\right|$. $\mathbf{i}$, Oscillating $\Delta I_{c}$, $\eta$ and fitted $\Delta I_{c}$ (red line). $\mathbf{j}$, Temperature-dependent $I_{c+}$, $I_{c-}$, $\Delta I_{c}$ and $\eta$ for device S2 at $B_{z}$ = 8.4 mT.
@@
The superconducting half-wave rectification is observed at $B_{z}$ = 8.4 mT and 10 mK in Fig. 1i, j. The ‘on’/‘off’ (superconducting/normal) states are switched by alternating current bias ($I_{b}=\pm$400 nA). Many cycles of these alternating measurements are conducted over a measurement time exceeding 1.5 hours, demonstrating the high stability of this rectification device, which is crucial for its potential applications. Intriguingly, the estimated switching power ($I_{b}^{2}R_{N}$) reaches the picowatt level (6.4 pW for device S2 with $R_{N}=40$ $\Omega$, 0.56 pW for device S1 in supplementary Fig. S1), which is four and eight orders of magnitude smaller than the field-free Josephson diode 25 and another bulk superconudcting diode 13, respectively. This power is also close to or even lower than the power of nanowire/nanoflake SDE systems 21; 19; 46; 47 and thin film SDE systems 15; 48; 18. The ultra-low switching power in Ta2Pd3Te5 Josephson diodes may be attributed to fewer edge supercurrent channels, making it a potential candidate for applications. Nonetheless, further technological advancements are required to optimize the efficiency, magnetic field conditions, and power consumption not only in this Josephson diode, but also in other superconducting diodes.
@@
To further study the origin of JDE, we measure the magnetic flux dependence of the JDE in the Ta2Pd3Te5 JJ. In Fig. 2a, $dV/dI$ as a function of $B_{z}$ and $I_{b}$ displays a SQUID pattern for device S1, suggesting the supercurrent interference of two channels in this JJ. The sawtooth-shaped $I_{c}$ pattern shows up, usually indicating the involvement of higher harmonic components due to non-negligible transmission channels 49, consistent with findings from the supplementary section I. The position-dependent supercurrent density distribution (Fig. 2b) extracted from the SQUID pattern 33 and the estimated enclosed area formed by two interference supercurrents (see supplementary section III) illustrate the supercurrent mainly originates from edge states of the JJ, confirming the reported edge state in Ta2Pd3Te5 42; 38; 40. In Fig. 2c, the $B_{z}$-dependent $I_{c+}$ and $\left|I_{c-}\right|$ extracted from Fig. 2a shows the asymmetric Josephson effect. Several features, such as the large deviation of $I_{c}$ from zero at half flux quantum in Fig. 2c and JDE in Fig. 2d, indicate the asymmetric edge supercurrent channels in this JJ, as detailed in methods. Moreover, $\Delta I_{c}=I_{c+}-\left|I_{c-}\right|$ oscillates with $B_{z}$ (orange curve in Fig. 2d), which can be called interfering JDE and is approximately described by the two-JJs model (red line in Fig. 2d, and see the detail in methods). This type of JDE has been theoretically predicted in Weyl semimetals with broken inversion symmetry and asymmetric helical edge states 5. Furthermore, the oscillating Josephson diode efficiency (cyan line in Fig. 2d) shows a maximal efficiency $\eta\approx$ 45 % at 6.4 mT. It is intriguing that $\eta$ can reach 10 % at $B_{z}$ = -0.5 mT. Also, the required field can be significantly reduced when the enclosed area of the SQUID loop is increased, which could be easily realized in the experimental setup 22; 21.
@@
Figure 3: Fractional Shapiro steps under microwave in both Ta2Pd3Te5 devices. $\mathbf{a}$, Differential resistance ($dV/dI$) as a function of voltage characteristics and flux quantum at microwave power = -32 dBm, $f$ = 5 GHz and 30 mK for device S1. The ±$\pm 1/2$th Shapiro steps are marked by black arrows. $\mathbf{b}$, $dV/dI$ versus $I_{b}$ at $B_{z}$ = 5.5 mT (3/2 $\phi_{0}$). Valleys characterized by half-integer steps are observed clearly. $\mathbf{c}$, Flux quantum and voltage dependence of $dV/dI$ at microwave power = 10 dBm, $f$ = 4.64 GHz and 10 mK for device S2. $\mathbf{d}$, Valleys characterized by the $-1/2$th Shapiro step at 6.7 mT (3/2 $\phi_{0}$). $\mathbf{e}$, Microwave power dependence of $dV/dI$ at 4.64 GHz and $B_{z}$ = 6.7 mT in device S2. $\mathbf{f}$, Microwave-power-dependent non-reciprocal critical current of the 0th Shapiro steps, which is extracted from $\mathbf{e}$.
@@
Having demonstrated the JDE in Ta2Pd3Te5 edge interferometer, we now show how to further enhance its efficiency. Interestingly, compared with device S1, the SQUID pattern of device S2 in Fig. 2f reveals two main periods, leading to a more pronounced asymmetry in the pattern (Fig. 2h, i) and an increased efficiency with $\eta_{max}$ = 73 % at 8.4 mT. The enhanced JDE can be explained by three JJs in parallel. The formation of the middle edge state in the JJ may stem from the ladder-like structure induced by imperfect exfoliation (as indicated by the green arrow in the inset of Fig. 2g and supplementary Fig. S3d), which is a common occurrence during the exfoliation process of 2D flakes 50 . These three supercurrent channels form two sets of interference patterns (corresponding red and green dashed loops in the inset of Fig. 2g), leading to the asymmetric SQUID. The $\Delta I_{c}$ can also be approximately fitted by the three-JJs model (red line in Fig. 2i, and see the detail in methods). The optimization of $\eta$ in device S2 (Fig. 2j), compared to device S1 in Fig. 2e, seems to be achieved by increasing the number of JJs in parallel, which is in line with the theoretically predicted method 9.
@@
Here, we examine the main mechanisms to determine which one best fits our data. (1) Magnetism mechanism 20 is initially excluded due to the absence of magnetism in the Ta2Pd3Te5 JJ. (2) Rashba-SOC mechanism is excluded as it typically requires in-plane magnetic fields, and Ising-SOC is also considered implausible because it is usually found in 2D superconductors 51. (3) JDE/SDE caused by finite-momentum Cooper pairing generally also needs in-plane magnetic fields 8; 14. The absence of obvious enhancement of the upper critical field under in-plane magnetic fields or the approximately butterfly interference pattern in our JJs 14 (supplementary Fig. S5) lead to the exclusion of finite-momentum Cooper pairing as a mechanism. (4) Nonlinear capacitance 6; 28 is absent in this system due to the symmetric superconducting electrodes and the lack of obvious hysteresis between switching current and retrapping current (Fig. 1h). (5) Self-inductance is also found to be too small to be considered (see supplementary section VIII). Therefore, the main origin of $B_{z}$-induced JDE in the Ta2Pd3Te5 edge interferometer is attributed to asymmetric edge supercurrents, as detailed in methods.
@@
As discussed above, besides asymmetric edge supercurrents and time-reversal symmetry breaking, the remaining higher harmonic in the CPR is further studied to explain the JDE in our JJs. The Shapiro resonances have been widely used to reveal not only the exotic symmetry of Cooper pairing 32 but also the higher harmonic or non-sinusoidal CPR 9; 52. The Shapiro steps appear in the $I$ $V$ curve at $V=V_{n}\equiv nhf/2e$ ($n$ is an integer) when the JJ is irradiated with the microwave. If the $n$th harmonic contributes to the CPR, $m/n$th Shapiro steps ($m$ is an integer) will be present 9; 52. In Fig. 3a, fractional Shapiro steps emerge at a microwave frequency of $f$ = 5 GHz, indicating the significant transparency of the JJ (supplementary section I). The ±$\pm 1/2$th Shapiro steps at half flux quantum $(n+1/2)\phi_{0}$ are marked by black arrows, which is also observed in other works 53; 24. The characteristic $dV/dI$ valleys formed by half-integer steps can be clearly observed at marked positions in Fig. 3b at $B_{z}$ = 5.5 mT (corresponding to 3/2 $\phi_{0}$). In addition, half-integer Shapiro steps are also observed in device S2 with flux-dependent measurements ($f$ = 4.64 GHz in Fig. 3c, d or 5.02 GHz in supplementary Fig. S7) and power-dependent measurements at $23\phi_{0}/$ 2 (Fig. 3e). These experiments indicate the existence of the second harmonic in the CPR, which is also supported by the rapid suppression observed in temperature-dependent $\Delta I_{c}$ compared with $I_{c}$ in device S1 (see Fig. 2e), because higher harmonics are quickly suppressed when temperature approaches $T_{c}$ 54. The rapid suppression of $\Delta I_{c}$ in device S2 (Fig. 2j) seems less pronounced, possibly because its three-JJs model enhances the diode effect, therefore relatively weakening such suppression. In addition to their applications in Josephson diodes, the JJs with higher harmonics also have potential applications in 0 - $\pi$ qubits, particularly if the first harmonic in the CPR can be effectively eliminated 55; 56; 57.

### v3

@@ Abstract
Edge states in topological systems have attracted great interest due to their robustness and linear dispersions. Here a superconducting-proximitized edge interferometer is engineered on a topological insulator Ta2Pd3Te5 with asymmetric edges to realize the interfering Josephson diode effect (JDE), which hosts many advantages, such as the high efficiency as much as 73% at tiny applied magnetic fields with an ultra-low switching power around picowatt. As an important element to induce such JDE, the second-order harmonic in the current-phase relation is also experimentally confirmed by half-integer Shapiro steps. The interfering JDE is also accompanied by the antisymmetric second harmonic transport, which indicates the corresponding asymmetry in the interferometer, as well as the polarity of JDE. This edge interferometer offers an effective method to enhance the performance of JDE, and boosts great potential applications for future superconducting quantum devices.
Introduction
@@
Figure 1: Mechanisms for SDE/JDE. $\mathbf{a}$, JDE in the asymmetric SQUID formed by two asymmetric JJs. $\mathbf{b}$, An asymmetric SQUID formed by asymmetric edge states in Ta2Pd3Te5 JJ and its device configuration. The upper and lower edges are the palladium and tantalum-tellurium atomic chain, respectively. The inset shows the photomicrograph of device S2, and the white scale bar corresponds to 1 $\mu$m. $\mathbf{c-f}$, Simulated CPR based on the minimal model. JDE only exists in $\mathbf{d}$ with different supercurrents, a non-zero magnetic field ($\phi/\phi_{0}\neq 0$) and higher harmonics in the CPR ($\alpha_{n}\neq 0$). The red and pink arrows represent the maximum($I_{c+}$) and minimum($I_{c-}$) current in the CPR, respectively. $\mathbf{g}$, Band structure of monolayer Ta2Pd3Te5. Edge states (blue and red lines) exist near the Fermi level and are marked in the corresponding edges of $\mathbf{b}$. $\mathbf{h}$, $\left|I_{b}\right|-\left|V\right|$ curves for positive and negative current sweep at $B_{z}$ = 8.4 mT. $\mathbf{i,j}$, Alternating switching between the superconducting and normal states at $B_{z}$ = 8.4 mT and 10 mK in device S2.
@@
Essentially, the appearance of JDE is related to the difference between maximum($I_{c+}$) and minimum($I_{c-}$) current in the CPR. The conventional CPR is an odd function of superconducting phase difference $\varphi$ between superconducting leads, entailing a zero supercurrent at $\varphi$ = 0, which shifts [meaning $I(\varphi=0)$neq 0] when both time-reversal and chiral symmetries are broken 30; 21; 31. To achieve JDE, other mechanisms to cause the CPR deviating from standard sinusoidal $I(\varphi)$ = $I_{c}$ $sin(\varphi)$ will also be introduced. The characteristics of CPR, which are usually inferred by the interference pattern of SQUIDs and the Shapiro steps due to ac Josephson effect, as well as the newly-discovered JDE with non-reciprocal critical current, have been widely harnessed to detect the anomalous superconductivity 32; 33, such as proximitized helical/chiral topological edge states. In turn, in a SQUID constructed by two asymmetric edge states of the topological insulator, JDE with an unconventional CPR could be easily realized with highly-tunable efficiency 5 and relatively easy fabrication (Fig. 1b). The relatively small number of edge supercurrent channels, in comparison to the case of bulk states transport, may contribute to the small critical current and power consumption 33; 34.
@@
More precisely, the interferometer formed by two asymmetric edge supercurrents (Fig. 1b) can be viewed as a SQUID with two different JJs, and induces JDE with requirements similar to the asymmetric SQUID 9 in Fig. 1a. According to the Souto $et$ $al$. 9, a minimal model with $n$ JJs concatenated in an interferometer array is applied to account for the origin of JDE, and the CPR only takes the first and second harmonic contribution into consideration. The current is written as $I(\varphi,\phi/\phi_{0})=\sum_{1}^{n=2}I_{n}sin(\varphi+2\pi(n-1)\phi/\phi_{0})+\alpha_{n}I_{n}sin(2\varphi+4\pi(n-1)\phi/\phi_{0})$ for two JJs in parallel, where $I_{n}$ ($\alpha_{n}I_{n}$) is the amplitude of the first (second) harmonic content for the $n$th JJ, $\alpha_{n}$ is current coefficient of the higher harmonic, $\phi$ is the magnetic flux and $\phi_{0}=h/2e$ is the magnetic flux quantum. In Fig. 1c-f, the simulations of the CPR using the aforementioned formula serves the purpose of illustrating three key requirements that contribute to the difference between $I_{c+}$ (red arrow) and $I_{c-}$ (pink arrow). First, an external magnetic field ($\phi/\phi_{0}\neq 0$) is necessary to break time-reversal symmetry and induce nonzero current at $\varphi$ = 0 in the CPR (Fig. 1d-f). Second, the supercurrent of each edge, including all the harmonic contributions, should be different (see Fig. 1d,e and detailed simulations in Supplementary Fig. 2). This condition can be realized by two interfering edge states with different dispersion 5. It can also be met, for example, by two superconducting quantum point contacts with different transmission coefficients 9, which may be linked to various factors such as different edge states, disorder situations, among others. All of these can cause the different supercurrent of edges in our devices, which is actually quite unique in topological materials. Finally, at least one edge channel should be transmissive so that the CPR acquires a higher harmonic ($\alpha_{n}\neq 0$) 9; 35 (see Fig. 1d, f).
Results
@@
In this work, highly efficient JDE, residing in the tilting supercurrent pattern, is reported in a Ta2Pd3Te5 edge interferometer at very small $B_{z}$ and is driven by low power consumption to achieve a highly stable rectification effect. It is also accompanied by fractional Shapiro steps, revealing the higher harmonic contributions of the non-negligible transmission channels. The choice of Ta2Pd3Te5 is based on the following reasons: (1) Ta2Pd3Te5 is a van der Waals material with quasi-one-dimensional (1D) chains 36 and can be easily mechanically exfoliated. (2) As a 2D topological insulator, where multiple layers could be viewed as simple stacking of monolayers due to very weak inter-layer coupling 37; 38, it hosts excitonic insulator states 39; 40; 41 and edge states observed by scanning tunnelling microscopy 42; 40 and transport measurements 38; 43. Notably, owing to the anisotropic bonding energy of 1D chains, the edges possess different atomic chains (Pd and Ta-Te atomic chains), resulting in asymmetric dispersion, as shown by the red and blue lines/arrows in Fig. 1b, g. (3) Superconductivity can be easily achieved in this system through doping or high-pressure techniques 44; 45; 46. Therefore, by proximity effect, the induced supercurrent (Supplementary Fig. 1) carried by edge states of Ta2Pd3Te5 can be used to realize an asymmetric SQUID with the interfering JDE.
Figure 2: Asymmetric Josephson effect and JDE in Ta2Pd3Te5 JJs. $\mathbf{a}$, SQUID pattern of device S1 at 15 mK. $\mathbf{b}$, Position-dependent supercurrent density distribution. Inset: A sketch of edge-supercurrents-formed JJ. $\mathbf{c}$, $B_{z}$-dependent $I_{c+}$ and $\left|I_{c-}\right|$. The switching current $I_{c+}$ and $I_{c-}$ are extracted by $\sim$ 15% of normal resistance. $\mathbf{d}$, Non-reciprocal critical current $\Delta I_{c}=I_{c+}-\left|I_{c-}\right|$, Josephson diode efficiency $\eta$ and fitted $\Delta I_{c}$ (black line) as a function of $B_{z}$. $\mathbf{e}$, Temperature-dependent $I_{c+}$, $I_{c-}$, $\Delta I_{c}$ and $\eta$ for device S1 at $B_{z}$ = 6.4 mT. The error bars primarily stem from the definition of the switching current, which is determined by 15% $\pm$ 5% drop of the normal resistance. $\mathbf{f}$, SQUID pattern of device S2 at 10 mK. $\mathbf{g}$, Supercurrent density distribution. The inset is a draft of the JJ with three edge supercurrents. $\mathbf{h}$, $B_{z}$ dependence of $I_{c+}$ and $\left|I_{c-}\right|$. $\mathbf{i}$, Oscillating $\Delta I_{c}$, $\eta$ and fitted $\Delta I_{c}$ (black line). $\mathbf{j}$, Temperature-dependent $I_{c+}$, $I_{c-}$, $\Delta I_{c}$ and $\eta$ for device S2 at $B_{z}$ = 8.4 mT. The error bars are obtained by the same method as in $\mathbf{e}$.
@@
The superconducting half-wave rectification is observed at $B_{z}$ = 8.4 mT and 10 mK in Fig. 1i, j. The ‘on’/‘off’ (superconducting/normal) states are switched by alternating current bias ($I_{b}=\pm$400 nA). Many cycles of these alternating measurements are conducted over a measurement time exceeding 1.5 hours, demonstrating the high stability of this rectification device, which is crucial for its potential applications. Intriguingly, the estimated switching power ($I_{b}^{2}R_{N}$) reaches the picowatt level (6.4 pW for device S2 with $R_{N}=40$ $\Omega$, 0.56 pW for device S1 in Supplementary Fig. 1), which is four and eight orders of magnitude smaller than the field-free Josephson diode 25 and another bulk superconudcting diode 13, respectively. This power is also close to or even lower than the power of nanowire/nanoflake SDE systems 21; 19; 47; 48 and thin film SDE systems 15; 49; 18. The ultra-low switching power in Ta2Pd3Te5 Josephson diodes may be attributed to fewer edge supercurrent channels, making it a potential candidate for applications. Nonetheless, further technological advancements are required to optimize the efficiency, magnetic field conditions, and power consumption not only in this Josephson diode, but also in other superconducting diodes.
@@
To further study the origin of JDE, we measure the magnetic flux dependence of the JDE in the Ta2Pd3Te5 JJ. In Fig. 2a, $dV/dI$ as a function of $B_{z}$ and $I_{b}$ displays a SQUID pattern for device S1, suggesting the supercurrent interference of two channels in this JJ. The sawtooth-shaped $I_{c}$ pattern shows up, usually indicating the involvement of higher harmonic components due to non-negligible transmission channels 50, consistent with findings from the Supplementary Note 1. The position-dependent supercurrent density distribution (Fig. 2b) extracted from the SQUID pattern 33 and the estimated enclosed area formed by two interference supercurrents (see Supplementary Note 3 and 9) illustrate the supercurrent mainly originates from edge states of the JJ, confirming the reported edge state in Ta2Pd3Te5 42; 38; 40. In Fig. 2c, the $B_{z}$-dependent $I_{c+}$ and $\left|I_{c-}\right|$ extracted from Fig. 2a shows the asymmetric Josephson effect. Several features, such as the large deviation of $I_{c}$ from zero at half flux quantum in Fig. 2c and JDE in Fig. 2d, indicate the asymmetric edge supercurrent channels in this JJ, as detailed in methods. Moreover, $\Delta I_{c}=I_{c+}-\left|I_{c-}\right|$ oscillates with $B_{z}$ (orange curve in Fig. 2d), which can be called interfering JDE and is approximately described by the two-JJs model (black line in Fig. 2d, and see the detail in methods). This type of JDE has been theoretically predicted in Weyl semimetals with broken inversion symmetry and asymmetric helical edge states 5. Furthermore, the oscillating Josephson diode efficiency (cyan line in Fig. 2d) shows a maximal efficiency $\eta\approx$ 45 % at 6.4 mT. It is intriguing that $\eta$ can reach 10 % at $B_{z}$ = -0.5 mT. Also, the required field can be significantly reduced when the enclosed area of the SQUID loop is increased, which could be easily realized in the experimental setup 22; 21.
@@
Figure 3: Fractional Shapiro steps under microwave in both Ta2Pd3Te5 devices. $\mathbf{a}$, Differential resistance ($dV/dI$) as a function of voltage characteristics and flux quantum at microwave power = -32 dBm, $f$ = 5 GHz and 30 mK for device S1. The ±$\pm 1/2$th Shapiro steps are marked by black arrows. $\mathbf{b}$, $dV/dI$ versus $I_{b}$ at $B_{z}$ = 5.5 mT (3/2 $\phi_{0}$). Valleys characterized by half-integer steps are observed clearly. $\mathbf{c}$, Flux quantum and voltage dependence of $dV/dI$ at microwave power = 10 dBm, $f$ = 4.64 GHz and 10 mK for device S2. $\mathbf{d}$, Valleys characterized by the $-1/2$th Shapiro step at 6.7 mT (3/2 $\phi_{0}$). $\mathbf{e}$, Microwave power dependence of $dV/dI$ at 4.64 GHz and $B_{z}$ = 6.7 mT in device S2. The $-1/2$th Shapiro step is marked by black arrows. $\mathbf{f}$, Microwave-power-dependent non-reciprocal critical current of the 0th Shapiro steps, which is extracted from $\mathbf{e}$. The switching currents here can be determined by peaks in the $dV/dI$ $I_{b}$ curve, as shown in Supplementary Fig. 8c, and the error bars primarily originate from the broadening of these peaks.
@@
Having demonstrated the JDE in Ta2Pd3Te5 edge interferometer, we now show how to further enhance its efficiency. Interestingly, compared with device S1, the SQUID pattern of device S2 in Fig. 2f reveals two main periods, leading to a more pronounced asymmetry in the pattern (Fig. 2h, i) and an increased efficiency with $\eta_{max}$ = 73 % at 8.4 mT. The enhanced JDE can be explained by three JJs in parallel. The formation of the middle edge state in the JJ may stem from the ladder-like structure induced by imperfect exfoliation (as indicated by the green arrow in the inset of Fig. 2g and supplementary Fig. 3d), which is a common occurrence during the exfoliation process of 2D flakes 51 . These three supercurrent channels form two sets of interference patterns (corresponding red and green dashed loops in the inset of Fig. 2g), leading to the asymmetric SQUID. The $\Delta I_{c}$ can also be approximately fitted by the three-JJs model (black line in Fig. 2i, and see the detail in methods). The optimization of $\eta$ in device S2 (Fig. 2j), compared to device S1 in Fig. 2e, seems to be achieved by increasing the number of JJs in parallel, which is in line with the theoretically predicted method 9.
@@
Here, we examine the main mechanisms to determine which one best fits our data. (1) Magnetism mechanism 20 is initially excluded due to the absence of magnetism in the Ta2Pd3Te5 JJ. (2) Rashba-spin orbital coupling (SOC) mechanism is excluded as it typically requires in-plane magnetic fields, and Ising-SOC is also considered implausible because it is usually found in 2D superconductors 52. (3) JDE/SDE caused by finite-momentum Cooper pairing generally also needs in-plane magnetic fields 8; 14. The absence of obvious enhancement of the upper critical field under in-plane magnetic fields or the approximately butterfly interference pattern in our JJs 14 (Supplementary Fig. 5) leads to the exclusion of finite-momentum Cooper pairing as a mechanism. (4) Nonlinear capacitance 6; 28 is absent in this system due to the symmetric superconducting electrodes and the lack of obvious hysteresis between switching current and retrapping current (Fig. 1h). (5) Self-inductance is also found to be too small to be considered (see Supplementary Note 8). Therefore, the main origin of $B_{z}$-induced JDE in the Ta2Pd3Te5 edge interferometer is attributed to asymmetric edge supercurrents, as detailed in methods.
@@
As discussed above, besides asymmetric edge supercurrents and time-reversal symmetry breaking, the remaining higher harmonic in the CPR is further studied to explain the JDE in our JJs. The Shapiro resonances have been widely used to reveal not only the exotic symmetry of Cooper pairing 32 but also the higher harmonic or non-sinusoidal CPR 9; 53. The Shapiro steps appear in the $I$ $V$ curve at $V=V_{n}\equiv nhf/2e$ ($n$ is an integer) when the JJ is irradiated with the microwave. If the $n$th harmonic contributes to the CPR, $m/n$th Shapiro steps ($m$ is an integer) will be present 9; 53. In Fig. 3a, fractional Shapiro steps emerge at a microwave frequency of $f$ = 5 GHz, indicating the significant transparency of the JJ (Supplementary Note 1). The ±$\pm 1/2$th Shapiro steps at half flux quantum $(n+1/2)\phi_{0}$ are marked by black arrows, which is also observed in other works 54; 24. The characteristic $dV/dI$ valleys formed by half-integer steps can be clearly observed at marked positions in Fig. 3b at $B_{z}$ = 5.5 mT (corresponding to 3/2 $\phi_{0}$). In addition, half-integer Shapiro steps are also observed in device S2 with flux-dependent measurements ($f$ = 4.64 GHz in Fig. 3c, d or 5.02 GHz in Supplementary Fig. 7) and power-dependent measurements at $23\phi_{0}/$ 2 (Fig. 3e). These experiments indicate the existence of the second harmonic in the CPR, which is also supported by the rapid suppression observed in temperature-dependent $\Delta I_{c}$ compared with $I_{c}$ in device S1 (see Fig. 2e), because higher harmonics are quickly suppressed when temperature approaches $T_{c}$ 55. The rapid suppression of $\Delta I_{c}$ in device S2 (Fig. 2j) seems less pronounced, possibly because its three-JJs model enhances the diode effect, therefore relatively weakening such suppression. In addition to their applications in Josephson diodes, the JJs with higher harmonics also have potential applications in 0 - $\pi$ qubits, particularly if the first harmonic in the CPR can be effectively eliminated 56; 57; 58.